# thanks: Corresponding author

### v1

@@ Abstract
where $P=(p+p^{\prime})/2$, $q=p^{\prime}-p$, and $t=q^{2}$, $a^{\{\mu}b^{\nu\}}=(a^{\mu}b^{\nu}+a^{\nu}b^{\mu})/2$. Here, $A(t)$, $B(t)$, and $D(t)$ represent the three fundamental GFFs of the nucleon. Consequently, investigating their properties is essential to unraveling the internal structure of the primary building blocks of matter. While the existing literature focuses predominantly on the GFFs $A(t)$ and $D(t)$, discussions regarding $B(t)$ remain relatively scarce [3, 4, 5]. Often, the form factor $J(t)=[A(t)+B(t)]/2$ is adopted in the literature. This observable is intimately linked to the nucleon’s spin structure; specifically, $B(0)$ corresponds to the anomalous gravitomagnetic moment (AGM). The equivalence principle (EP) requires that $B(0)=0$ [3], a result proven in light-front QCD [6] and supported by lattice simulations [7, 8] and model-based computations [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Furthermore, recent lattice QCD simulations suggest that the nucleon GFF $B_{N}(t)$ remains consistent with zero even at finite $t$, within current simulation uncertainties [7, 8]. Although some dispersion relation analyses based on lattice data suggest a non-vanishing $B_{N}(t)$ [19, 20, 21], the extracted values are sufficiently small that they are often neglected in practical applications, such as predicting near-threshold $J/\psi$ photoproduction [22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Regardless, the persistent smallness of $B_{N}(t)$ across different scales warrants a formal explanation.
@@
The suppression of $B_{N}(t)$ can be further understood through the fundamental symmetries governing graviton interactions. Within the Dyson-Schwinger equations (DSEs) framework, $B_{N}(t)$ is found to be vanishingly small, strictly obeying the constraint $B_{N}(0)=0$ [17]. This result is primarily driven by the symmetry of the graviton-quark vertex, which is rigorously enforced by Ward-Takahashi identities derived from the EP [34]. Complementary insights are provided by holographic QCD, which models hadron-graviton interactions within a 5D warped geometry. In this framework, GFFs are computed via the scattering of a bulk graviton off a nucleon field in Anti-de Sitter (AdS) space. Notably, in models employing minimal gravitational coupling, $B_{N}(t)$ vanishes identically across all momentum transfers [35]. This result is fundamentally rooted in the fact that the minimal interaction vertex in the 5D bulk conserves helicity in the massless limit. Since the form factor $B(t)$ is the gravitational analog of the Pauli form factor $F_{2}(t)$ and specifically corresponds to a helicity-flip transition, the absence of a helicity-flipping mechanism in the minimal coupling scheme leads to its vanishing. Consequently, any non-zero value for $B_{N}(t)$ would require the introduction of non-minimal coupling terms, such as higher-derivative Pauli-type interactions in the 5D action. The empirical observation that $B_{N}(t)$ is nearly zero suggests that the nucleon’s interaction with gravity is dominated by these minimal, helicity-conserving holographic processes.
A closely related approach is light-front holographic QCD, which establishes a mapping between the holographic description in 5D AdS space and light-front QCD [36, 37] (see Ref. [38] for a review). Within this framework, the AGM $B_{N}(0)$ vanishes identically for each individual Fock sector, a result fundamentally ensured by the underlying light-front dynamics. The GFF $B_{N}(t)$ at finite $t$ also vanishes in this approach, as directly implied by the holographic correspondence. However, it is important to note that for theories other than QCD, the GFF $B(t)$ is not required to be zero. By performing a comparative analysis of the light-front wave functions (LFWFs) across different theoretical frameworks, we can gain deeper insight into the specific underlying dynamics that suppress $B_{N}(t)$ for the nucleon. This suggests that the smallness of this observable is not merely a general property of field theories, but a specific consequence of the nucleon’s unique internal structure.
On the light-front side, the nucleon emerges as a bound state consisting of an active quark $q$ and a scalar diquark $D$. Within this representation, the GFF $B_{N}(t)=B_{q}(t)+B_{D}(t)$ is extracted from the spin-flip matrix element $T^{++}_{\uparrow\downarrow}$. The specific wave function representations for the quark and diquark contributions are given by [38]:
@@
In light-front holographic QCD, the wave functions adopt a remarkably simple form where the explicit dependence on the longitudinal momentum fraction $x$ effectively drops out:
@@
As a result, the total current $\mathcal{B}^{(0)}(z,-Q^{2})$ vanishes identically. This leads to a vanishing GFF $B_{N}(t)$ for all finite $t$, providing a microscopic light-front derivation that is perfectly consistent with the predictions of holographic QCD in the 5D bulk.
A central assumption in the above derivation in light-front holographic QCD is that the wave functions depend exclusively on the holographic variable $\zeta_{\perp}$. However, as demonstrated in Refs. [39, 40], a longitudinal wave function $X(x)$ is generally required to account for longitudinal dynamics. Incorporating this component, the modified wave functions are expressed as:
@@
Figure 1: The longitudinal wave function $X_{\pm}(x)=X(x)$, the corresponding suppression factor $X^{2}(x)-X^{2}(1-x)$, and the integrand of Eq. (11) at $zQ=1$ with two sets of mass parameters: (Top) $m_{q}=46$ MeV and $m_{D}=140$ MeV; (Bottom) $m_{q}=300$ MeV, $m_{D}=600$ MeV. The former are the values adopted in light-front holographic QCD and the latter are typical values adopted in constituent quark model.
In light-front holographic QCD, both the active quark and the diquark cluster are massless. Finite quark masses cause the deviation of the longitudinal wave function from unity, becoming asymmetric. A standard ansatz for the longitudinal wave function in this framework is [40]:
@@
where $N$ is a normalization constant, $\kappa=0.54$ GeV represents the confining strength, and $m_{q}=46$ MeV and $m_{D}=140$ MeV are the quark and diquark masses, respectively. Despite the lack of perfect symmetry in this ansatz, the integral remains significantly suppressed due to the antisymmetric factor, as illustrated in Fig. 1. For comparison, Fig. 1 also presents a second set of longitudinal wave functions utilizing masses typical of the constituent quark model ($m_{q}=300$ MeV, $m_{D}=600$ MeV). While these constituent-like masses increase the magnitude of the integrand, the total integral remains small because of internal cancellations within the antisymmetric weighting. Figure 2 illustrates the numerical results for the GFF $B_{N}(t)$, obtained by incorporating the longitudinal wave function from Eq. (12) into the light-front holographic QCD framework. The result explicitly demonstrates the suppression of $B_{N}(t)$ across the physical range of momentum transfer.
Figure 2: The gravitational form factor $B_{N}(t)$ of the nucleon obtained from light-front holographic QCD (LFHQCD) incorporating longitudinal dynamics. The theoretical prediction is shown in comparison with recent lattice QCD results from Ref. [8].
In Fig. 3, we compare the integrands from different partial waves. They are obtained by multiplying the wave function $X(x)$ given in Eq. (12) by the Legendre polynomial $P_{n}(2x-1)$, where $n=0,1,2$ for S-, P-, D-waves, with a proper normalization. As shown in Fig. 3, the longitudinal wave functions and corresponding integrands for S-, P-, and D-waves reveal that the S-wave component undergoes the most dramatic suppression. Consequently, within light-front holographic QCD, the observed smallness of the GFF $B_{N}(t)$ at finite $t$ serves as a signature of the nucleon’s dominant S-wave character.
@@
In the context of GPDs, the smallness of $B_{N}(t)$ is largely attributed to a flavor cancellation between the $u$ and $d$ quark sectors. Meanwhile, in holographic QCD, the vanishing of $B_{N}(t)$ under minimal coupling emerges as a direct consequence of helicity conservation in the 5D bulk. However, neither interpretation is entirely satisfactory: the former relies on fine-tuning within the model parameters, while the latter is strictly limited to the minimal coupling scheme – conflicting with lattice QCD results and other models where $B_{N}(t)$ is non-zero.
By bridging these descriptions through light-front holographic QCD, we demonstrate that this suppression is analytically tied to an antisymmetric factor inherent in the longitudinal dynamics. At the forward limit ($t=0$), the suppression is exact as required by the EP. At finite momentum transfer $t$, the integral governing $B_{N}(t)$ vanishes identically if the longitudinal wave functions are symmetric and remain heavily suppressed even when realistic mass scales are introduced. This suggests that the near-zero value of $B_{N}(t)$ is a distinctive signature of the nucleon’s dominant S-wave character. While our analysis is grounded in light-front holographic QCD, the underlying mechanism presented here is remarkably robust: it relies on fundamental symmetry properties and the relativistic structure of the nucleon rather than the specific fine-tuning of model parameters. Consequently, this mechanism likely remains applicable to a broader class of relativistic constituent models where $B_{N}(t)$ is small but non-vanishing.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC) under Grant Nos. 12375081, 12247103, and by the Chinese Academy of Sciences under Grant No. YSBR-101. X.-h. Cao is supported by the NSFC under Grant No. 125B2111. C. M. is supported by new faculty start up funding the Institute of Modern Physics, Chinese Academy of Sciences, Grants No. E129952YR0.

### v2

@@ Abstract
where $P=(p+p^{\prime})/2$, $q=p^{\prime}-p$, and $t=q^{2}$, $a^{\{\mu}b^{\nu\}}=(a^{\mu}b^{\nu}+a^{\nu}b^{\mu})/2$. Here, $A(t)$, $B(t)$, and $D(t)$ represent the three fundamental GFFs of the nucleon. Often, the form factor $J(t)=[A(t)+B(t)]/2$ is adopted in the literature. Consequently, investigating their properties is essential to unraveling the internal structure of the primary building blocks of matter.
While the existing literature focuses predominantly on the GFFs $A(t)$ and $D(t)$, discussions regarding $B(t)$ remain relatively scarce [3, 4, 5]. This observable is intimately linked to the nucleon’s spin structure; specifically, $B(0)$ corresponds to the anomalous gravitomagnetic moment (AGM). The equivalence principle (EP) requires that $B(0)=0$ [3], a result proven in light-front QCD [6] and supported by lattice simulations [7, 8] and model-based computations [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. The vanishing of $B(0)$ is consistent with the tensor meson dominance in $\pi N$ scattering [20, 21]. Furthermore, recent lattice QCD simulations suggest that the nucleon GFF $B_{N}(t)$ remains consistent with zero even at finite $t$, within current simulation uncertainties [7, 8]. Although some dispersion relation analyses based on lattice data suggest a non-vanishing $B_{N}(t)$ as shown in Fig. 1 [22, 23, 24], the extracted values are sufficiently small that they are often neglected in practical applications, such as predicting near-threshold $J/\psi$ photoproduction [25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Regardless, the persistent smallness of $B_{N}(t)$ across different scales warrants a formal explanation.
Figure 1: The gravitational form factor $B_{N}(t)$ of the nucleon obtained from lattice QCD (dots, [8]), dispersion relation and VMD analyses based on the lattice data (bands, [22, 23, 24]), as well as two theoretical models, LFHQCD (solid line, see texts) and DSEs (dot-dash lines, [18]).
@@
The suppression of $B_{N}(t)$ can be further understood through the fundamental symmetries governing graviton interactions. Within the Dyson-Schwinger equations (DSEs) framework, $B_{N}(t)$ is found to be vanishingly small, strictly obeying the constraint $B_{N}(0)=0$ [18]. This result is primarily driven by the symmetry of the graviton-quark vertex, which is rigorously enforced by Ward-Takahashi identities derived from the EP [37]. Complementary insights are provided by holographic QCD, which models hadron-graviton interactions within a 5D warped geometry [38, 39, 25, 26, 16, 40, 41, 42, 43, 44, 45, 46, 47, 48]. In this framework, GFFs are computed via the scattering of a bulk graviton off a nucleon field in Anti-de Sitter (AdS) space. Notably, in models employing minimal gravitational coupling, $B_{N}(t)$ vanishes identically across all momentum transfers [39]. This result is fundamentally rooted in the fact that the minimal interaction vertex in the 5D bulk conserves helicity in the massless limit. Since the form factor $B(t)$ is the gravitational analog of the Pauli form factor $F_{2}(t)$ and specifically corresponds to a helicity-flip transition, the absence of a helicity-flipping mechanism in the minimal coupling scheme leads to its vanishing. Consequently, any non-zero value for $B_{N}(t)$ would require the introduction of non-minimal coupling terms, such as higher-derivative Pauli-type interactions in the 5D action. The empirical observation that $B_{N}(t)$ is nearly zero suggests that the nucleon’s interaction with gravity is dominated by these minimal, helicity-conserving holographic processes.
A closely related approach is light-front holographic QCD (LFHQCD), which establishes a mapping between the holographic description in 5D AdS space and light-front QCD [49, 50] (see Ref. [51] for a review). Within this framework, the AGM $B_{N}(0)$ vanishes identically for each individual Fock sector, a result fundamentally ensured by the underlying light-front dynamics. The GFF $B_{N}(t)$ at finite $t$ also vanishes in this approach, as directly implied by the holographic correspondence. However, it is important to note that for theories other than QCD, the GFF $B(t)$ is not required to be zero [6, 5]. By performing a comparative analysis of the light-front wave functions (LFWFs) across different theoretical frameworks, we can gain deeper insight into the specific underlying dynamics that suppress $B_{N}(t)$ for the nucleon. This suggests that the smallness of this observable is not merely a general property of field theories, but a specific consequence of the nucleon’s unique internal structure.
On the light-front side, the nucleon emerges as a bound state consisting of an active quark $q$ and a scalar diquark $D$. Within this representation, the GFF $B_{N}(t)=B_{q}(t)+B_{D}(t)$ is extracted from the spin-flip matrix element $T^{++}_{\uparrow\downarrow}$. The specific wave function representations for the quark and diquark contributions are given by [6, 52]:
@@
In LFHQCD, the wave functions adopt a remarkably simple form where the explicit dependence on the longitudinal momentum fraction $x$ effectively drops out [51]:
@@
As a result, the total current $\mathcal{B}^{(0)}(z,-Q^{2})$ vanishes identically. This leads to a vanishing GFF $B_{N}(t)$ for all finite $t$, providing a microscopic light-front derivation that is perfectly consistent with the predictions of holographic QCD in the 5D bulk [39].
A central assumption in the above derivation in LFHQCD is that the wave functions depend exclusively on the holographic variable $\zeta_{\perp}$. However, as demonstrated in Refs. [53, 54], a longitudinal wave function $X(x)$ is generally required to account for longitudinal dynamics. Incorporating this component, the modified wave functions are expressed as:
@@
Figure 2: The longitudinal wave function $X_{\pm}(x)=X(x)$, the corresponding suppression factor $X^{2}(x)-X^{2}(1-x)$, and the integrand of Eq. (11) at $zQ=1$ with two sets of mass parameters: (Top) $m_{q}=46$ MeV and $m_{D}=140$ MeV; (Bottom) $m_{q}=300$ MeV, $m_{D}=600$ MeV. The former are the values adopted in LFHQCD and the latter are typical values adopted in constituent quark model.
In LFHQCD, both the active quark and the diquark cluster are massless. Finite quark masses cause the deviation of the longitudinal wave function from unity, becoming asymmetric. A standard ansatz for the longitudinal wave function in this framework is [54]:
@@
where $N$ is a normalization constant, $\kappa=0.54$ GeV represents the confining strength, and $m_{q}=46$ MeV and $m_{D}=140$ MeV are the quark and diquark masses, respectively. Despite the lack of perfect symmetry in this ansatz, the integral remains significantly suppressed due to the antisymmetric factor, as illustrated in Fig. 2. For comparison, Fig. 2 also presents a second set of longitudinal wave functions utilizing masses typical of the constituent quark model ($m_{q}=300$ MeV, $m_{D}=600$ MeV). While these constituent-like masses increase the magnitude of the integrand, the total integral remains small because of internal cancellations within the antisymmetric weighting. Figure 1 illustrates the numerical results for the GFF $B_{N}(t)$, obtained by incorporating the longitudinal wave function from Eq. (12) into the LFHQCD framework. The result explicitly demonstrates the suppression of $B_{N}(t)$ across the physical range of momentum transfer.
@@
In Fig. 3, we compare the integrands from different partial waves. They are obtained by multiplying the wave function $X(x)$ given in Eq. (12) by the Legendre polynomial $P_{n}(2x-1)$, where $n=0,1,2$ for S-, P-, D-waves, with a proper normalization. As shown in Fig. 3, the longitudinal wave functions and corresponding integrands for S-, P-, and D-waves reveal that the S-wave component undergoes the most dramatic suppression. Consequently, within LFHQCD, the observed smallness of the GFF $B_{N}(t)$ at finite $t$ serves as a signature of the nucleon’s dominant S-wave character. Figure 4 compares the GFF $B(t)$ across various longitudinal partial waves. Notably, the S-wave contribution is significantly suppressed in comparison to the higher-order partial waves.
Figure 4: Gravitational form factor $B(t)$ calculated from various partial wave contributions to the longitudinal wave function. These wave functions are derived using an invariant mass ansatz, where the $n$-th partial wave is represented by the Legendre polynomial $P_{n}(2x-1)$.
@@
In the context of GPDs, the smallness of $B_{N}(t)$ is largely attributed to a flavor cancellation between the $u$ and $d$ quark sectors. Meanwhile, in holographic QCD, the vanishing of $B_{N}(t)$ under minimal coupling emerges as a direct consequence of helicity conservation in the 5D bulk. However, neither interpretation is entirely satisfactory: the former relies on fine-tuning within the model parameters, while the latter is strictly limited to the minimal coupling scheme — conflicting with lattice QCD results and other models where $B_{N}(t)$ is non-zero.
By bridging these descriptions through LFHQCD, we demonstrate that this suppression is analytically tied to an antisymmetric factor inherent in the longitudinal dynamics. At the forward limit ($t=0$), the suppression is exact as required by the EP. At finite momentum transfer $t$, the integral governing $B_{N}(t)$ vanishes identically if the longitudinal wave functions are symmetric and remain heavily suppressed even when realistic mass scales are introduced. This suggests that the near-zero value of $B_{N}(t)$ is a distinctive signature of the nucleon’s dominant S-wave character. While our analysis is grounded in LFHQCD, the underlying mechanism presented here is remarkably robust: it relies on fundamental symmetry properties and the relativistic structure of the nucleon rather than the specific fine-tuning of model parameters. Consequently, this mechanism likely remains applicable to a broader class of relativistic constituent models where $B_{N}(t)$ is small but non-vanishing.
@@ Acknowledgements
The authors thank C.D. Roberts and Xionghui Cao for providing the gravitational form factor data calculated via Dyson-Schwinger equations (DSEs) and dispersion relations (DR), respectively. This work is supported by the National Natural Science Foundation of China (NSFC) under Grant Nos. 12375081, 12247103, and by the Chinese Academy of Sciences under Grant No. YSBR-101. X.-h. Cao is supported by the NSFC under Grant No. 125B2111. C. M. is supported by new faculty start up funding the Institute of Modern Physics, Chinese Academy of Sciences, Grants No. E129952YR0.